# Vision Transformers on the Edge: A Comprehensive Survey of Model Compression and Acceleration Strategies

### v1

@@ 4.3.3 Throughput
Table 14 illustrates the throughput of the accelerating techniques. We observe significant variations in throughput, influenced by factors such as hardware architecture, precision, and optimization strategies. Zhang et al. (2330.2 GOPs) achieve the highest throughput, leveraging an FPGA-based implementation with optimized parallel execution. Auto-ViT-Acc (1181.5 GOPs) and M³ViT (1217.4 GOPs) also report high throughput, suggesting effective hardware utilization. However, several studies, such as ViTCoD, do not report the throughput. Additionally, EQ-ViT and VAQF balance throughput with power efficiency, offering a more energy-efficient alternative.
@@ 5 Challenges and Future Directions of ViT on Edge Devices
ViT models are computation-intensive, and their deployment on resource-constrained edge devices has been a big challenge. However, with the advancement of edge AI, this is now changing, and the efficient and cost-effective implementation of ViT models is possible directly on edge hardware. This increases accessibility for end users and reduces reliance on cloud infrastructure, which lowers latency, improves privacy, and reduces operational costs. However, some areas, such as real-world scenarios and software-hardware co-design still need to be explored for ViT on edge devices. In this section, we will discuss the current challenges and future opportunities of ViT on edge devices.
5.1 Software-Hardware Co-design
@@
The lightweight ViT model and compression techniques should be considered the hardware architecture. The SW-HW co-design can reduce the current dilemma between model and hardware architectures. Additionally, different edge hardware platforms (e.g., CPUs, GPUs, and FPGAs) have varying capabilities in handling precision, memory bandwidth, and computational efficiency. Often, accelerators support a uniform bit-width tensor, and this distinct bit-width precision needs zero padding, incurring inefficient memory usage. It is so hard to optimize the ViT for each type of hardware. Leveraging hardware-aware compression techniques can improve the efficiency of edge deployment. Frameworks such as DNNWeaver [128], VAQF [119], M3ViT [19] have been developed for different hardware platforms FPGA, GPU accelerators for efficient edge inference. However, most of the current framework can not handle the sparsity caused by model compression. Therefore, the advancement of reconfigurability of software-hardware co-design for handling sparsity can be a future problem to solve.
5.2 Utlizing NAS for Inference
NAS is currently used to find the optimized model. The limitations are the computational load and time required for the NAS algorithms, which are high. Consequently, improving NAS algorithms to obtain optimized ViT models targeting the edge can be a future research opportunity. Moreover, frameworks like HAQ [59] and APQ [129] utilized NAS for automatically generating pruning quantization strategies through reinforcement learning or evolutionary search methods. However, this work is highly customized for specific hardware (e.g., HAQ for FPGA). This problem arises because different edge hardware platforms or neural accelerators have distinct properties and processing capabilities. ProxylessNAS [130] is one of the works that can find a model to fit the hardware but is limited to only CNN models. Therefore, utilizing NAS for searching hardware-independent optimization techniques for ViT can be the future direction to explore.
5.3 Acclerators to Handle Sparsity
Traditional processors such as GPU, CPU, or even FPGA cannot handle the space, irregular tensor. For example, although mixed-precision quantization techniques for ViT have been developed, their deployment on edge devices remains limited due to the inefficiency of current hardware architectures and accelerators in handling mixed-precision formats [131]. Therefore, optimized accelerators to handle mixed precision format on edge devices still need to be explored. In addition, combining multiple compression techniques for optimum hardware performance is a promising research direction. Current hardware accelerators are not inherently designed to process sparse tensors efficiently, as they require fetching zero values from memory to processing elements (PEs). Thus, specialized techniques are needed to optimize the storage and computation of nonzero values in ViT.
@@ 5.4 Automated Edge Aware Model Compression
Most of the current model compression techniques require manual adjustment of hyperparameters such as quantization bit width, pruning ratio, or layer-wise sparsity. Compression hyperparameters must be adjusted automatically or adaptively within the resource budget with minimum degradation of accuracy. From our observation, few works explore adjusting the compression parameters automatically. SparseViT [132] is one of few works that effectively reduced computation by targeting less-important regions with dynamically chosen pruning ratios in the images for the ViT model, achieving significant latency reductions. As a result, developing a hardware-efficient automaticity compression technique can be an interesting research domain in the future.
@@ 5.5 Developing Benchmarks
Proper benchmark standards to evaluate the performance of the edge devices are important. The different stages of the model to deploying edge, including compression and accelerators, require a universal and comprehensive set of metrics to compare different proposed solutions. However, benchmarking datasets and models from the system perspective are limited. For example, most of the compression techniques for ViT were evaluated on the ImageNet-1k [51] dataset for classification tasks and COCO-2017 [91] datasets for object detection tasks. However, expanding that knowledge for different real-world application areas is still limited due to the lack of dataset and model benchmarks. Thus, more benchmarking datasets and ViT models for evaluating the proposed system/framework need to be developed for different application areas, such as medical imaging and autonomous driving. In addition, making one compression technique universal for different CV tasks is challenging. Making task-independent universal compression techniques can be an interesting research domain in the future.
@@ 5.6 Real-world Case Studies
The deployment of ViT on edge devices has gained significant progress in recent years. However, most of the compression techniques, frameworks, and accelerators are limited to evaluation in an academic environment. For instance, there are numerous ViT model on medical imaging datasets for different CV tasks such as image classification [133, 134, 135, 136, 137], segmentation [138, 10, 139, 140, 141], object detection [142, 143, 144, 145, 146]. However, few studies have evaluated compression and accelerator techniques on those ViT-based medical imaging models. Exploring ViT on edge for medical imaging can be challenging because of the unique nature of the data (3D ultrasounds or MRIs).
Maintaining accuracy, latency, and precision is critical in real-world applications, particularly in critical fields like medical imaging. In medical imaging, a significant challenge to the major compressing of the ViT models is maintaining the spatial resolution and feature details since even small degradations in accuracy will affect the diagnostic outcome significantly [147]. Compression techniques can reduce excessive feature abstraction, potentially discarding vital low-level details essential for accurate diagnoses. Furthermore, transfer learning in medical imaging adds another layer of complexity—determining which pre-trained layers to retain or modify without losing critical learned representations is a significant challenge [148, 149]. Therefore, it is an open research direction to achieve a balance between model efficiency and diagnostic reliability for ViT models on edge for real-world scenarios such as healthcare applications.
@@ 5.7 Seamless Model-to-Edge Integration
The conversion from a trained model into a hardware-compatible version for inference requires extensive time, cost, and, most importantly, manual in each step. Additionally, there is a high knowledge gap between the research community. For example, training new models requires extensive software knowledge, while compressing and developing accelerating strategies require deep hardware knowledge. It is challenging to find an expert in both directions in the research community. These difficulties create a significant research gap in developing tools that automatically map the models on hardware. However, FPGA can overcome some limitations with the ability to reconfigure new operations and modules. However, the available tools are insufficient for the automatic mapping of models and are more limited for ViT.
@@ 6 Conclusion
With the increasing adoption of ViTs in computer vision, optimizing their efficiency for edge deployment has become a key research focus. This survey comprehensively analyzes ViT model compression and hardware-aware acceleration techniques, exploring techniques such as pruning, quantization, knowledge distillation, and SW-HW co-design. By categorizing these advancements and evaluating their impact across edge platforms, we highlight the trade-offs between accuracy, resource utilization, and energy efficiency in real-world applications. Our analysis indicates that while SOTA ViT compression and acceleration techniques effectively reduce computational overhead and improve inference speed, challenges such as hardware adaptability, memory bottlenecks, and optimal compression strategies remain unexplored. Additionally, We discuss the potential future directions, such as utilizing NAS to find hardware-aware optimization parameters, sparsity-aware accelerators, and efficient cross-platform SW-HW co-design frameworks.
Overall, the domain of optimizing ViT on edge devices remains an evolving field, presenting both challenges and opportunities. Continued advancements in co-optimized software-hardware solutions will pave the way for more efficient and deployable ViTs.
@@ 7 Acknowledgement
This work was partly supported by the U.S. National Science Foundation under Grants CNS-2245729 and Michigan Space Grant Consortium 80NSSC20M0124.

### v2

@@ 4.3.3 Throughput
Table 17 illustrates the throughput of the accelerating techniques. We observe significant variations in throughput, influenced by factors such as hardware architecture, precision, and optimization strategies. Zhang et al. (2330.2 GOPs) [133] achieve the highest throughput, leveraging an FPGA-based implementation with optimized parallel execution. Auto-ViT-Acc (1181.5 GOPs) and M³ViT (1217.4 GOPs) also report high throughput, suggesting effective hardware utilization. However, several studies, such as ViTCoD, do not report the throughput. Additionally, EQ-ViT and VAQF balance throughput with power efficiency, offering a more energy-efficient alternative.
@@ 5 Challenges and Future Directions of ViT on Edge Devices
ViT models are computation-intensive, and their deployment on resource-constrained edge devices has been a big challenge. However, with the advancement of edge AI, this is now changing, and the efficient and cost-effective implementation of ViT models is possible directly on edge hardware. This increases accessibility for end users and reduces reliance on cloud infrastructure, which lowers latency, improves privacy, and reduces operational costs. However, some areas, such as real-world scenarios and SW-HW co-design still need to be explored for ViT on edge devices. In this section, we will discuss the current challenges and future opportunities of ViT on edge devices.
5.1 Advancement in SW-HW Co-design
@@
The SW-HW co-design can reduce the current dilemma between model and hardware architectures. Additionally, different edge hardware platforms (e.g., CPUs, GPUs, and FPGAs) have varying capabilities in handling precision, memory bandwidth, and computational efficiency. Often, accelerators support a uniform bit-width tensor, and this distinct bit-width precision needs zero padding, incurring inefficient memory usage. It is so hard to optimize the ViT for each type of hardware. Leveraging hardware-aware compression techniques can improve the efficiency of edge deployment. Frameworks such as DNNWeaver [136], VAQF [122], M3ViT [19] have been developed for different hardware platforms like FPGA, GPU accelerators for efficient edge inference. However, most of the current frameworks can not handle the sparsity caused by model compression. Therefore, the advancement of reconfigurability of SW-HW co-design for handling sparsity can be a future problem to solve.
Additionally, future work should focus on establishing collaborative SW-HW design pipelines that integrate compression-aware modeling with hardware-specific execution constraints. For example, model compression in software could be dynamically adapted based on hardware feedback or user-specific hardware requirements. This requires tighter integration between models and hardware profiling tools. Further, hardware-agnostic intermediate representations and unified compiler toolchains could help bridge the gap between diverse ViT models and heterogeneous edge devices.
5.2 Utilizing NAS for Inference and Compression Parameter Search
NAS is widely used to discover efficient model structures. However, its application to ViTs on edge devices remains limited due to high computational costs and strong coupling to specific hardware platforms. Frameworks such as HAQ [85] and APQ [137] utilized NAS to automate pruning and quantization strategies through reinforcement learning or evolutionary algorithms. However, these works are highly customized for specific hardware (e.g., HAQ for FPGA). ProxylessNAS [138] is a notable effort in this direction that introduces hardware-aware search and demonstrates strong real-time latency prediction, but it is limited to CNN architectures. Extending such approaches to ViT and its hierarchical, attention-based structure remains an open challenge.
Future research should extend NAS to simultaneously search for model-independent compression parameters, such as quantization bit-widths, pruning ratios, and distillation strategies, based on hardware-aware objectives (e.g., latency, memory usage, energy efficiency). Reinforcement learning can be utilized to efficiently guide the search process based on hardware requirements. Additionally, exploring NAS to FPGA custom design, which requires deep hardware expertise, might be an interesting and challenging research area for the efficient deployment of ViT.
5.3 Hardware Accelerators to Handle Sparsity and Mixed Precision
Traditional processors such as GPUs, CPUs, or even FPGAs struggle with executing irregular tensor operations efficiently. Current accelerators are not inherently designed to process sparse tensors, as they often fetch zero values from memory to processing elements (PEs), incurring unnecessary latency and energy overhead. Another challenge is the limited support for mixed-precision execution. Although recent techniques have shown that mixed-precision quantization can preserve ViT accuracy while reducing model size, most edge hardware requires uniform bit-widths across operations [139], leading to compatibility issues or bit padding and making it hard to deploy on edge devices.
To address these limitations, future research should focus on developing sparsity-aware accelerator architectures capable of dynamically skipping zero computations, using methods such as compressed sparse row (CSR) encoding, zero-skipping logic [140], or sparsity-driven scheduling techniques [141, 142]. Similarly, designing bit-level reconfigurable compute units could enable native support for mixed-precision execution without introducing significant hardware redundancy. Thus, specialized techniques are needed to optimize the storage and computation of nonzero values in ViT.
Additionally, emerging math-based accelerating techniques such as tensor-train (TT) decomposition offer a promising direction for accelerating ViT models. TT-decomposition can significantly reduce parameter count and computational complexity by representing high-dimensional weight matrices in a compact, low-rank format. Recent studies, including TT-ViT [143] and TT-MLP [144], demonstrate its potential for transformer and MLP compression, while hardware-aware implementations [145] further highlight its relevance for deployment. Although integration into edge-specific ViT deployments remains limited, TT-decomposition represents a promising future research avenue for ViT acceleration, especially on FPGA due to its reconfigurable property.
@@ 5.4 Automated Edge Aware Model Compression
Most of the current model compression techniques require manual adjustment of hyperparameters such as quantization bit width, pruning ratio, or layer-wise sparsity. Compression hyperparameters must be adjusted automatically or adaptively within the resource budget with minimum degradation of accuracy. From our observation, few works explore adjusting the compression parameters automatically. SparseViT [146] is one of the works that effectively reduced computation by targeting less-important regions with dynamically chosen pruning ratios in the images for the ViT model, achieving significant latency reductions. As a result, developing a hardware-efficient automaticity compression technique can be an interesting research domain in the future.
@@ 5.5 Developing Benchmarks
Proper benchmark standards to evaluate the performance of the edge devices are essential. The different stages of the model for deploying edge, including compression and accelerators, require a universal and comprehensive set of metrics to compare different proposed solutions. However, benchmarking datasets and models from the system perspective is limited. For example, most of the compression techniques for ViT were evaluated on the ImageNet-1k [45] dataset for classification tasks and the COCO-2017 [96] datasets for object detection tasks. However, expanding that knowledge to different real-world application areas is still limited due to the lack of datasets and model benchmarks. Thus, more benchmarking datasets and ViT models for evaluating the proposed system/framework need to be developed for different application areas, such as medical imaging and autonomous driving. In addition, making one compression technique universal for different CV tasks is challenging. Making task-independent universal compression techniques can be an interesting research domain in the future.
@@ 5.6 Real-world Case Studies
The deployment of ViT on edge devices has gained significant progress in recent years. However, most of the compression techniques, frameworks, and accelerators are limited to evaluation in an academic environment. For instance, there are numerous ViT models on medical imaging datasets for different CV tasks such as image classification [147, 148, 149], segmentation [150, 10, 151, 152, 153], object detection [154, 155, 156, 157], even more specialized tasks like vessel-level image quality assessment [158]. However, few studies have evaluated compression and accelerator techniques on those ViT-based medical imaging models. Exploring ViT on edge for medical imaging can be challenging because of the unique nature of the data (3D ultrasounds or MRIs). Additionally, recent advancements have extended the application of ViTs beyond high-level vision tasks to encompass more specialized low-level tasks. Notable examples include flow-guided video inpainting using global-local aggregation [159], and adaptive spatiotemporal graph transformer networks for action and image quality assessment [160]. However, the deployment of such models on edge devices remains largely unexplored. Investigating the applicability of ViTs to these low-level, temporally complex tasks in resource-constrained environments presents a promising direction for future research.
Maintaining accuracy, latency, and precision is critical in real-world applications, particularly in critical fields like medical imaging. In medical imaging, a significant challenge during compression of the ViT models is maintaining the spatial resolution and feature details since even small degradations in accuracy will affect the diagnostic outcome significantly [161]. Compression techniques can reduce excessive feature abstraction, potentially discarding vital low-level details essential for accurate diagnoses. Furthermore, transfer learning in medical imaging adds another layer of complexity—determining which pre-trained layers to retain or modify without losing critical learned representations is a significant challenge [162, 163]. Therefore, it is an open research direction to achieve a balance between model efficiency and diagnostic reliability for ViT models on edge for real-world scenarios such as healthcare applications.
@@ 5.7 Seamless Model-to-Edge Integration
The conversion from a trained model into a hardware-compatible version for inference requires extensive time, cost, and, most importantly, manual labor in each step. Additionally, there is a high knowledge gap between the research community. For example, training new models requires extensive software knowledge, while compressing and developing accelerating strategies require deep hardware knowledge. It is challenging to find an expert in both directions in the research community. These difficulties create a significant research gap in developing tools that automatically map the models to hardware. However, FPGA can overcome some limitations with the ability to reconfigure new operations and modules. However, the available tools are insufficient for the automatic mapping of models and are more limited for ViT.
@@ 6 Conclusion
In this survey, we comprehensively analyze model compression and hardware acceleration techniques on ViT models, spanning techniques ranging from pruning, quantization, and KD, software tools, to SW-HW co-design. Our analysis indicates that while SOTA ViT compression and acceleration techniques effectively reduce computational overhead and improve inference speed, challenges such as hardware adaptability, memory bottlenecks, and optimal compression strategies remain unexplored. Additionally, we discuss the potential future directions, such as utilizing NAS to find hardware-aware optimization parameters, sparsity-aware accelerators, and efficient cross-platform SW-HW co-design frameworks.
Overall, optimizing ViT for edge devices is an evolving field, marked by both challenges and opportunities. This survey provides a comprehensive overview of the current landscape and aims to inspire continued innovation in this crucial area.
@@ 7 Acknowledgement
This work was partly supported by the U.S. National Science Foundation under Grants CNS-2245729 and Michigan Space Grant Consortium 80NSSC20M0124. The authors would also like to thank Yunge Li for the continuous discussions and insightful feedback that helped shape the direction of this study.